# Can a calibration metric be both testable and actionable?

### v1

@@ Appendix B Decision-theoretic guarantees: proofs and extensions
$R_{\mathrm{bd}}(f;\tau)-R_{\mathrm{bd}}(h\circ f;\tau)$ (45)
$=\tau\mathbb{E}[(1-Y)({{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}-{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}})]+(1-\tau)\mathbb{E}[Y({{\mathbf{1}_{\left\{f(X)<\tau\right\}}}}-{{\mathbf{1}_{\left\{h\circ f(X)<\tau\right\}}}})]$ (46)
@@
$\begin{split}R_{\mathrm{bd}}(f;\tau)-R_{\mathrm{bd}}(h\circ f;\tau)=&\tau\mathbb{E}[(1-Y){{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)<\tau\right\}}}}]\\ &-\tau\mathbb{E}[(1-Y){{\mathbf{1}_{\left\{f(X)<\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}}]\\ &+(1-\tau)\mathbb{E}[Y{{\mathbf{1}_{\left\{f(X)<\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}}]\\ &-(1-\tau)\mathbb{E}[Y{{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)<\tau\right\}}}}]\end{split}$ (47)
We achieve the desired result by distributing $(1-\tau)$ and $(1-Y)$ and collecting like terms. ∎
@@
    $\leq\mathbb{E}[(\mathbb{E}[Y\mid f(X)]-f(X)){{\mathbf{1}_{\left\{f<\tau\right\}}}}{{\mathbf{1}_{\left\{\mathbb{E}[Y\mid f(X)]\geq\tau\right\}}}}]+\mathbb{E}[(f(X)-\mathbb{E}[Y\mid f(X)]){{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}{{\mathbf{1}_{\left\{\mathbb{E}[Y\mid f(X)]<\tau\right\}}}}]$
@@
$R_{\mathrm{bd}}(f;\tau)-R_{\mathrm{bd}}(h\circ f;\tau)$ (48)
$=\mathbb{E}[(Y-\tau){{\mathbf{1}_{\left\{f(X)<\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}}]+\mathbb{E}[(\tau-Y){{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)<\tau\right\}}}}]$ (49)
$\leq\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f<\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}}]+\mathbb{E}[(f(X)-Y){{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}{{\mathbf{1}_{\left\{h(f)<\tau\right\}}}}]$ (50)
$=\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)<\tau\right\}}}}{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]+\mathbb{E}[(f(X)-Y){{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}{{\mathbf{1}_{\left\{f(X)\in I^{\prime}\right\}}}}]$ (51)
$\leq 2\Delta_{\operatorname{Cutoff}}(f)$ (52)
@@
We now generalize our results for $\ell_{\mathrm{bd}}$ to a broader class of loss functions, still in the context of $Y\in\{0,1\}$. Like Hu and Wu (2024) and Kleinberg et al. (2023), we consider a loss function :{0,1}×[$\ell:\{0,1\}\times[0,1]\rightarrow\mathbb{R}$ that follows the very minimal constraint that it be a proper scoring rule:
$\mathbb{E}_{Y\sim\mathrm{Bernoulli}(p)}[\ell(Y,p)]\leq\mathbb{E}_{Y\sim\mathrm{Bernoulli}(q)}[\ell(Y,p)]$ (53)
@@
Suppose $Y\in\{0,1\}$. Suppose :{0,1}×[$\ell:\{0,1\}\times[0,1]\rightarrow\mathbb{R}$ is a proper scoring rule that also satisfies the regularity conditions listed in Gneiting and Raftery (2007, Theorem 3). If the measure $\nu$ used to define the Schervish representation is a probability measure, then
$\mathbb{E}[\ell(Y,f(X))]-\inf_{h}\mathbb{E}[\ell(Y,h\circ f(X))]\leq\Delta_{\operatorname{ECE}}(f)$ (54)
@@
$\mathbb{E}[\ell(Y,f(X))]-\inf_{h:\textnormal{ monotone}}\mathbb{E}[\ell(Y,h\circ f(X))]\leq 2\Delta_{\operatorname{Cutoff}}(f)$ (55)
@@
    $\mathbb{E}[\ell(Y,f(X))]-\mathbb{E}[\ell(Y,h\circ f(X))]=\mathbb{E}\left[\mathbb{E}_{\tau\sim\nu}[\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}},\tau)-\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}},\tau)]\right]$ (56)
@@
    $\mathbb{E}\left[\mathbb{E}_{\tau\sim\nu}[\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}},\tau)-\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}},\tau)]\right]$ (57)
    $=\mathbb{E}_{\tau\sim\nu}\left[\mathbb{E}[\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}},\tau)-\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}},\tau)]\right]$ (58)
    $\leq\sup_{\tau\in[0,1]}\mathbb{E}[\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}},\tau)-\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}},\tau)]$ (59)
@@ B.4 Cutoff Calibration Error upper bounds the monotone risk gap of a sign-testing loss
$R_{\mathrm{st}}(Y,f)=\mathbb{E}[(Y-Y^{*}){{\mathbf{1}_{\left\{f(X)\leq Y^{*}\right\}}}}{{\mathbf{1}_{\left\{Y>Y^{*}\right\}}}}]+\mathbb{E}[(Y^{*}-Y){{\mathbf{1}_{\left\{f(X)>Y^{*}\right\}}}}{{\mathbf{1}_{\left\{Y\leq Y^{*}\right\}}}}]$ (60)
@@
$\begin{split}R_{\mathrm{st}}(f)-R_{\mathrm{st}}(h\circ f)=&\mathbb{E}[(Y-Y^{*}){{\mathbf{1}_{\left\{f(X)\leq Y^{*}\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)>Y^{*}\right\}}}}]\\ &+\mathbb{E}[(Y^{*}-Y){{\mathbf{1}_{\left\{f(X)>Y^{*}\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)\leq Y^{*}\right\}}}}]\end{split}$ (61)
@@
$R_{\mathrm{st}}(f)-\inf_{h\textnormal{ monotone}}R_{\mathrm{st}}(h(f))\leq 2\Delta_{\operatorname{Cutoff}}(f)$ (62)
@@
$R_{\mathrm{st}}(f)-\inf_{h}R_{\mathrm{st}}(h(f))\leq\Delta_{\operatorname{ECE}}(f)$ (63)
@@
The fact that we recover the same theoretical guarantees as before is somewhat intuitive, given that the Bayes decision rule here has the same form as before: ${{\mathbf{1}_{\left\{\mathbb{E}[Y\mid f]\geq Y^{*}\right\}}}}$
@@ B.5 Implementation details of the experiment
$\sup_{w\in\mathcal{L}_{1}}\frac{1}{N}\sum_{j=1}^{N}w(f(X_{j}))(\mathbb{E}[Y_{j}\mid X_{j}]-f(X_{j})).$ (64)
@@
$\sup_{I:\textnormal{ interval}}\frac{1}{N}\sum_{j=1}^{N}(\mathbb{E}[Y_{j}\mid X_{j}]-f(X_{j})){{\mathbf{1}_{\left\{f(X_{j})\in I\right\}}}}.$ (65)
@@
$\frac{1}{N}\sum_{j=1}^{N}\ell_{\mathrm{bd}}(\mathbb{E}[Y_{j}\mid X_{j}],{{\mathbf{1}_{\left\{f(X_{j})\geq\tau\right\}}}};\tau).$ (67)
@@
$\frac{1}{N}\sum_{j=1}^{N}\ell_{\mathrm{bd}}(\mathbb{E}[Y_{j}\mid X_{j}],{{\mathbf{1}_{\left\{\mathbb{E}[Y_{j}\mid X_{j}]\geq\tau\right\}}}};\tau).$ (68)
@@
$\begin{split}\min\biggl{(}&\min_{\tau^{\prime}\in\{0,f(X_{1}),\dots,f(X_{N}),1\}}\frac{1}{N}\sum_{j=1}^{N}\ell_{\mathrm{bd}}(\mathbb{E}[Y_{j}\mid X_{j}],{{\mathbf{1}_{\left\{f(X_{j})\geq\tau^{\prime}\right\}}}};\tau),\\ &\min_{\tau^{\prime}\in\{0,f(X_{1}),\dots,f(X_{N}),1\}}\frac{1}{N}\sum_{j=1}^{N}\ell_{\mathrm{bd}}(\mathbb{E}[Y_{j}\mid X_{j}],{{\mathbf{1}_{\left\{f(X_{j})\leq\tau^{\prime}\right\}}}};\tau)\biggr{)}.\end{split}$ (69)
@@ Appendix C Testability: proofs
$\mathcal{A}:(\mathcal{X}\times[0,1])^{n}\rightarrow\{\text{ measurable functions }f:\mathcal{X}\rightarrow[0,1]\}$ (70)
be any procedure that inputs a data set $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$, and returns a fitted function $f_{n}=\mathcal{A}\big{(}(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\big{)}$. Let $(X_{1},Y_{1}),\dots,(X_{N},Y_{N})\overset{iid}{\sim}P$, let $f_{n}=\mathcal{A}\big{(}(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\big{)}$ be trained on the first $n$ data points, and let $\widehat{P}_{N-n}=\frac{1}{N-n}\sum_{i=n+1}^{N}\delta_{(X_{i},Y_{i})}$ be the empirical distribution of the remaining $N-n$ data points. then
@@
$\Delta_{\operatorname{ECE}}^{\widehat{P}_{N-n}}(f)\geq\frac{1}{N-n}\sum_{i=n+1}^{N}|Y_{i}-f(X_{i})|\cdot{{\mathbf{1}_{\left\{f(X_{i})\neq f(X_{j}),\leavevmode\nobreak\ \forall j\in\{n+1,\dots,N\}\backslash\{i\}\right\}}}}.$
@@
$\mathbb{E}\left[\Delta_{\operatorname{ECE}}^{\widehat{P}_{N-n}}(f)\right]=\frac{1}{N-n}\sum_{i=n+1}^{N}\mathbb{E}\left[|Y_{i}-f(X_{i})|\cdot{{\mathbf{1}_{\left\{f(X_{i})\neq f(X_{j}),\leavevmode\nobreak\ \forall j\in\{n+1,\dots,N\}\backslash\{i\}\right\}}}}\right],$
@@
$\textnormal{Var}(Y\mid X=x)=\mathbb{E}[|Y-\mathbb{E}[Y\mid X]|^{2}\mid X=x]=\inf_{t\in[0,1]}\mathbb{E}[|Y-t|^{2}\mid X=x]\leq\inf_{t\in[0,1]}\mathbb{E}[|Y-t|\mid X=x],$
@@
Note that we must have $<N-n$ many values $t\in[0,1]$ such that $\mathbb{P}_{P}(f(X)=t)>\frac{1}{N-n}$, so
@@
If $S_{\gamma}(f,P)\geq N-n$, then, we must have $\sup_{t_{1},\dots,t_{N-n}}\mathbb{P}_{P}(f(X)\in\{t_{1},\dots,t_{N-n-1}\})<1-\gamma$ and so $\mathbb{P}_{P}(p(X)\leq\frac{1}{N-n})\geq\gamma$. Therefore,
If Sγ(f,P)Nn, then $\textnormal{If $S_{\gamma}(f,P)\geq N-n$, then $\mathbb{E}\left[\big{(}Y_{n+1}-\mathbb{E}[Y_{n+1}\mid X_{n+1}]\big{)}^{2}\cdot{{\mathbf{1}_{\left\{p(X_{n+1})\leq\frac{1}{N-n}\right\}}}}\right]\geq\sigma^{2}_{\gamma}(P)$},$
@@
    where $i_{1},\dots,i_{n}\overset{iid}{\sim}\mathrm{Uniform}([N])$. Now, with probability $(1-\frac{1}{N})\cdot\ldots\cdot(1-\frac{n-1}{N})\geq 1-\frac{n(n-1)}{2N}$, the indices $i_{1},\dots,i_{n}$ are all distinct—i.e., equivalently, these indices have been sampled without replacement. Therefore,
@@
    $\mathbb{E}_{P}[w(f(X))\cdot(Y-f(X))]\leq\mathbb{E}_{P}[w(f(X))\cdot(Y-f(X))]+2d_{TV}(P,Q),$
@@
    $\leq\sup_{w:[0,1]\to[-1,1]}\mathbb{E}_{Q}[w(f(X))\cdot(Y-f(X))]$
    $=\Delta_{\operatorname{ECE}}^{Q}(f).$

### v2

@@ Appendix B Decision-theoretic guarantees: proofs and extensions
    $R_{\mathrm{bd}}(f;\tau)-R_{\mathrm{bd}}(h\circ f;\tau)$
    $=\tau\mathbb{E}[(1-Y)({{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}-{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}})]+(1-\tau)\mathbb{E}[Y({{\mathbf{1}_{\left\{f(X)<\tau\right\}}}}-{{\mathbf{1}_{\left\{h\circ f(X)<\tau\right\}}}})].$
@@
    $\begin{split}R_{\mathrm{bd}}(f;\tau)-R_{\mathrm{bd}}(h\circ f;\tau)=&\tau\mathbb{E}[(1-Y){{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)<\tau\right\}}}}]\\ &-\tau\mathbb{E}[(1-Y){{\mathbf{1}_{\left\{f(X)<\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}}]\\ &+(1-\tau)\mathbb{E}[Y{{\mathbf{1}_{\left\{f(X)<\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}}]\\ &-(1-\tau)\mathbb{E}[Y{{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)<\tau\right\}}}}].\end{split}$
    We achieve the desired result by distributing $(1-\tau)$ and (1Y)(1-Y) and collecting like terms. ∎
@@
    $\leq\mathbb{E}[(\mathbb{E}[Y\mid f(X)]-f(X)){{\mathbf{1}_{\left\{f(X)<\tau\right\}}}}{{\mathbf{1}_{\left\{\mathbb{E}[Y\mid f(X)]\geq\tau\right\}}}}]+\mathbb{E}[(f(X)-\mathbb{E}[Y\mid f(X)]){{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}{{\mathbf{1}_{\left\{\mathbb{E}[Y\mid f(X)]<\tau\right\}}}}]$
@@
    $R_{\mathrm{bd}}(f;\tau)-R_{\mathrm{bd}}(h\circ f;\tau)$
    $=\mathbb{E}[(Y-\tau){{\mathbf{1}_{\left\{f(X)<\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}}]+\mathbb{E}[(\tau-Y){{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)<\tau\right\}}}}]$
    $\leq\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)<\tau\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}}]+\mathbb{E}[(f(X)-Y){{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}{{\mathbf{1}_{\left\{h(f)<\tau\right\}}}}]$
    $=\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)<\tau\right\}}}}{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]+\mathbb{E}[(f(X)-Y){{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}}{{\mathbf{1}_{\left\{f(X)\in I^{\prime}\right\}}}}]$
    $\leq 2\Delta_{\operatorname{Cutoff}}(f).$
@@
We now generalize our results for $\ell_{\mathrm{bd}}$ to a broader class of loss functions, still in the context of $Y\in\{0,1\}$. Like Hu and Wu (2024) and Kleinberg et al. (2023), we consider a loss function $\ell:\{0,1\}\times[0,1]\rightarrow\mathbb{R}$ that follows the very minimal constraint that it be a proper scoring rule:
$\mathbb{E}_{Y\sim\mathrm{Bernoulli}(p)}[\ell(Y,p)]\leq\mathbb{E}_{Y\sim\mathrm{Bernoulli}(q)}[\ell(Y,p)].$
@@
Suppose $Y\in\{0,1\}$. Suppose $\ell:\{0,1\}\times[0,1]\rightarrow\mathbb{R}$ is a proper scoring rule that also satisfies the regularity conditions listed in Gneiting and Raftery (2007, Theorem 3). If the measure $\nu$ used to define the Schervish representation is a probability measure, then
$\mathbb{E}[\ell(Y,f(X))]-\inf_{h}\mathbb{E}[\ell(Y,h\circ f(X))]\leq\Delta_{\operatorname{ECE}}(f)$
@@
$\mathbb{E}[\ell(Y,f(X))]-\inf_{h:\textnormal{ monotone}}\mathbb{E}[\ell(Y,h\circ f(X))]\leq 2\Delta_{\operatorname{Cutoff}}(f).$
@@
    $\mathbb{E}[\ell(Y,f(X))]-\mathbb{E}[\ell(Y,h\circ f(X))]=\mathbb{E}\left[\mathbb{E}_{\tau\sim\nu}[\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}},\tau)-\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}},\tau)]\right]$
@@
    $\mathbb{E}\left[\mathbb{E}_{\tau\sim\nu}[\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}},\tau)-\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}},\tau)]\right]$
    $=\mathbb{E}_{\tau\sim\nu}\left[\mathbb{E}[\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}},\tau)-\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}},\tau)]\right]$
    $\leq\sup_{\tau\in[0,1]}\mathbb{E}[\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{f(X)\geq\tau\right\}}}},\tau)-\ell_{\mathrm{bd}}(Y,{{\mathbf{1}_{\left\{h\circ f(X)\geq\tau\right\}}}},\tau)]$
@@ B.4 Cutoff Calibration Error upper bounds the monotone risk gap of a sign-testing loss
$R_{\mathrm{st}}(f;Y^{*})=\mathbb{E}[(Y-Y^{*}){{\mathbf{1}_{\left\{f(X)\leq Y^{*}\right\}}}}{{\mathbf{1}_{\left\{Y>Y^{*}\right\}}}}]+\mathbb{E}[(Y^{*}-Y){{\mathbf{1}_{\left\{f(X)>Y^{*}\right\}}}}{{\mathbf{1}_{\left\{Y\leq Y^{*}\right\}}}}].$
@@
$\begin{split}R_{\mathrm{st}}(f;Y^{*})-R_{\mathrm{st}}(h\circ f;Y^{*})=&\mathbb{E}[(Y-Y^{*}){{\mathbf{1}_{\left\{f(X)\leq Y^{*}\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)>Y^{*}\right\}}}}]\\ &+\mathbb{E}[(Y^{*}-Y){{\mathbf{1}_{\left\{f(X)>Y^{*}\right\}}}}{{\mathbf{1}_{\left\{h\circ f(X)\leq Y^{*}\right\}}}}].\end{split}$
@@
For any $Y^{*}\in[0,1]$,
$R_{\mathrm{st}}(f;Y^{*})-\inf_{h\textnormal{ monotone}}R_{\mathrm{st}}(h\circ f;Y^{*})\leq 2\Delta_{\operatorname{Cutoff}}(f)$
@@
$R_{\mathrm{st}}(f;Y^{*})-\inf_{h}R_{\mathrm{st}}(h\circ f;Y^{*})\leq\Delta_{\operatorname{ECE}}(f).$
@@
The fact that we recover the same theoretical guarantees as before is somewhat intuitive, given that the Bayes decision rule here has the same form as before: $\mathbf{1}_{\left\{\mathbb{E}[Y\mid f]\geq Y^{*}\right\}}$}}.
@@ B.5 Implementation details of the experiment
In this section, we give details for the implementation of the experiment in Section 3.3.
@@
$\sup_{w\in\mathcal{L}_{1}}\frac{1}{N}\sum_{j=1}^{N}w(f(X_{j}))(\mathbb{E}[Y_{j}\mid X_{j}]-f(X_{j})).$
@@
$\sup_{I:\textnormal{ interval}}\frac{1}{N}\sum_{j=1}^{N}(\mathbb{E}[Y_{j}\mid X_{j}]-f(X_{j})){{\mathbf{1}_{\left\{f(X_{j})\in I\right\}}}}.$
@@
$\frac{1}{N}\sum_{j=1}^{N}\ell_{\mathrm{bd}}(\mathbb{E}[Y_{j}\mid X_{j}],{{\mathbf{1}_{\left\{f(X_{j})\geq\tau\right\}}}};\tau).$
@@
$\frac{1}{N}\sum_{j=1}^{N}\ell_{\mathrm{bd}}(\mathbb{E}[Y_{j}\mid X_{j}],{{\mathbf{1}_{\left\{\mathbb{E}[Y_{j}\mid X_{j}]\geq\tau\right\}}}};\tau).$
@@
$\begin{split}\min\biggl{(}&\min_{\tau^{\prime}\in\{0,f(X_{1}),\dots,f(X_{N}),1\}}\frac{1}{N}\sum_{j=1}^{N}\ell_{\mathrm{bd}}(\mathbb{E}[Y_{j}\mid X_{j}],{{\mathbf{1}_{\left\{f(X_{j})\geq\tau^{\prime}\right\}}}};\tau),\\ &\min_{\tau^{\prime}\in\{0,f(X_{1}),\dots,f(X_{N}),1\}}\frac{1}{N}\sum_{j=1}^{N}\ell_{\mathrm{bd}}(\mathbb{E}[Y_{j}\mid X_{j}],{{\mathbf{1}_{\left\{f(X_{j})\leq\tau^{\prime}\right\}}}};\tau)\biggr{)}.\end{split}$
@@ Appendix C Testability: proofs
$\mathcal{A}:(\mathcal{X}\times[0,1])^{n}\rightarrow\{\text{ measurable functions }f:\mathcal{X}\rightarrow[0,1]\}$
be any procedure that inputs a data set $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$, and returns a fitted function $f_{n}=\mathcal{A}\big{(}(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\big{)}$. Let $(X_{1},Y_{1}),\dots,(X_{N},Y_{N})\overset{iid}{\sim}P$, let $f_{n}=\mathcal{A}\big{(}(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\big{)}$ be trained on the first $n$ data points, and let $\widehat{P}_{N-n}=\frac{1}{N-n}\sum_{i=n+1}^{N}\delta_{(X_{i},Y_{i})}$ be the empirical distribution of the remaining NnN-n data points. then
@@
[…]
@@
[…]
@@
$\textnormal{Var}(Y\mid X=x)$ $=\mathbb{E}[|Y-\mathbb{E}[Y\mid X]|^{2}\mid X=x]$
$=\inf_{t\in[0,1]}\mathbb{E}[|Y-t|^{2}\mid X=x]\leq\inf_{t\in[0,1]}\mathbb{E}[|Y-t|\mid X=x],$
@@
Note that we must have <Nn<N-n many values $t\in[0,1]$ such that $\mathbb{P}_{P}(f(X)=t)>\frac{1}{N-n}$, so
@@
If $S_{\gamma}(f,P)\geq N-n$, then, we must have $\sup_{t_{1},\dots,t_{N-n-1}}\mathbb{P}_{P}(f(X)\in\{t_{1},\dots,t_{N-n-1}\})<1-\gamma$ and so $\mathbb{P}_{P}(p(X)\leq\frac{1}{N-n})\geq\gamma$. Therefore,
$\textnormal{If $S_{\gamma}(f,P)\geq N-n$, then $\mathbb{E}\left[\textnormal{Var}(Y_{n+1}|X_{n+1})\cdot{{\mathbf{1}_{\left\{p(X_{n+1})\leq\frac{1}{N-n}\right\}}}}\right]\geq\sigma^{2}_{\gamma}(P)$},$
@@
    where $i_{1},\dots,i_{n}\overset{iid}{\sim}\mathrm{Uniform}([N])$. Now, with probability $(1-\frac{1}{N})\cdot\ldots\cdot(1-\frac{n-1}{N})\geq 1-\frac{n(n-1)}{2N}$, the indices $i_{1},\dots,i_{n}$ are all distinct—i.e., equivalently, these indices have been sampled without replacement (WOR). Therefore,
@@
    $\mathbb{E}_{P}[w(f(X))\cdot(Y-f(X))]\leq\mathbb{E}_{Q}[w(f(X))\cdot(Y-f(X))]+2d_{TV}(P,Q),$
@@
    $\leq\sup_{w:[0,1]\to[-1,1]}\mathbb{E}_{Q}[w(f(X))\cdot(Y-f(X))]+2d_{TV}(P,Q)$
    $=\Delta_{\operatorname{ECE}}^{Q}(f)+2d_{TV}(P,Q).$